Can someone help me solve this problem?
Solve the system by graphing both equations on the same axes. Give the ordered pair for the point of intersection.
x+2y=4
6y=-3x+12

Answer by Alan3354(69443)

Solve the system by graphing both equations on the same axes. Give the ordered pair for the point of intersection.
x+2y=4
6y=-3x+12 --> x+2y=4
----------
It's the same equation.
-----------
It's a straight line, so you need 2 points.
x+2y=4
x = 4 - 2y
Pick 2 values for y and find x.
Plot the 2 points.
Draw a line thru them.